### v1

# CDFlow: Building Invertible Layers with Circulant and Diagonal Matrices

Abstract
Normalizing flows are deep generative models that achieve efficient likelihood estimation and sampling through invertible transformations. A key challenge is designing linear layers that enhance expressiveness while enabling efficient computation of the Jacobian determinant and inverse. In this work, we introduce a novel invertible linear layer based on the product of circulant and diagonal matrices. This decomposition provides a parameter- and computation-efficient formulation, reducing the parameter complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(mn)$ by using $m$ diagonal matrices together with m1m-1 circulant matrices, while approximating arbitrary linear transformations. Furthermore, leveraging the Fast Fourier Transform (FFT), our method reduces the time complexity of matrix inversion from $\mathcal{O}(n^{3})$ to $\mathcal{O}(mn\log n)$ and matrix log-determinant from $\mathcal{O}(n^{3})$ to $\mathcal{O}(mn)$, where $n$ is the input dimension. Building upon this, we introduce a novel normalizing flow model called Circulant-Diagonal Flow (CDFlow). Empirical results demonstrate that CDFlow excels in density estimation for natural image datasets and effectively models data with inherent periodicity. In terms of computational efficiency, our method speeds up the matrix inverse and log-determinant computations by $1.17\times$ and $4.31\times$, respectively, compared to the general dense matrix, when the number of channels is set to 96.
1 Introduction
Deep generative models have gained significant attention due to their potential applications across various domains, including natural language processing (Dong et al., 2024), image restoration and generation (Yao et al., 2023), robot learning (Chao et al., 2024), and autonomous driving (Guan et al., 2024). Among these models, normalizing flows hold a distinctive position because of their ability to provide accurate likelihood estimation and efficient sampling.
Flow-based models enable the probabilistic modeling of complex distributions by transforming them into simpler ones through a series of invertible mappings. Their uniqueness lies in the strict requirements for the construction of these functions: on one hand, the functions must be invertible; on the other hand, the determinant of the corresponding Jacobian matrix should be easy to compute. However, these constraints may limit the expressive power of the models.
To address this challenge, NICE (Dinh et al., 2014) introduced additive coupling layers, which were later extended by RealNVP (Dinh et al., 2016) to affine coupling layers. These coupling layers have since become a fundamental component of most flow-based models. However, the design of these non-linear layers, which only transform a subset of variables, requires coordination with linear layers to effectively capture dependencies across arbitrary dimensions. As a result, recent efforts have focused on developing specialized linear layer designs, ranging from fixed channel-disrupting operations (Dinh et al., 2014, 2016) to $1\times 1$ convolutions (Kingma and Dhariwal, 2018) and even $d\times d$ convolutions (Hoogeboom et al., 2019) with learnable parameters. However, dense convolutions can result in significant memory overhead as a result of the large weight matrices required for high-dimensional inputs. One approach to address this is the construction of reversible butterfly layers (Meng et al., 2022) using butterfly matrices (Dao et al., 2019), which help maintain the model’s expressiveness while reducing the number of parameters.
Structured matrices and Fourier transform acceleration have been demonstrated in numerous studies (Cheng et al., 2015; Ding et al., 2017; Karami et al., 2019; Jensen et al., 2021) to be highly effective in reducing the number of parameters and computational complexity of neural networks. Research by (Huhtanen and Perämäki, 2015) demonstrates that any matrix $\mathbf{M}\in\mathbb{R}^{n\times n}$ can be represented as the alternating product of a cyclic matrix and a diagonal matrix, requiring no more than 2n12n-1 factors. This factorization structure has demonstrated versatility and effectiveness in domains such as deep neural networks  (Moczulski et al., 2015; Araujo et al., 2019) and parameter-efficient fine-tuning  (Ding et al., 2025). However, these studies have not focused on model reversibility.
Inspired by previous work, we propose a novel normalizing flow model called Circulant-Diagonal Flow (CDFlow), which is built on invertible linear layers constructed using circulant and diagonal matrices. This construction preserves the expressive power of the model while significantly reducing storage demands, as only the diagonal and circulant elements need to be stored instead of the full $n\times n$ weight matrix. In addition, the distinct properties of circulant and diagonal matrices enable the use of the Fast Fourier Transform (FFT) to enhance computation efficiency (Ding et al., 2025). This facilitates faster calculations for both the determinant of Jacobian matrices and the inversion of weight matrices. Overall, the main contributions of our work can be summarized as following:
  • We propose CDFlow, a new class of flow-based generative models. The linear layer in CDFlow is designed to have a storage complexity that scales linearly with the input dimension, as it only requires storing a collection of circulant and diagonal vectors — namely, the diagonal elements of diagonal matrices and the eigenvalue vectors of circulant matrices. In practice, the linear layer weights can be constructed using a small number of circulant and diagonal matrices, enabling efficient transformations across both spatial and channel dimensions.
  • We leverage the unique properties of circulant and diagonal matrices to significantly reduce computational complexity. A key property is that the determinant of any diagonalizable matrix equals the product of its eigenvalues. Accordingly, by performing an $\mathcal{O}(n\log n)$ FFT once during initialization to precompute the eigenvalues of circulant matrices, we enable $\mathcal{O}(n)$ log-determinant evaluation at runtime. Additionally, the complexity of matrix inversion is reduced to $\mathcal{O}(n\log n)$.
  • We demonstrate through image generation experiments that our CDFlow performs effectively while also offering superior computational efficiency in terms of both log-determinant computation and matrix inverse.
2 Background
2.1 Flow-based Models
The normalizing flow model maps the complex data distribution $p_{X}(x)$ to a known simple probability distribution $p_{Z}(z)$ via an invertible transformation $f_{\theta}:x\to z$ for data generation with $\theta$ being the model parameter. The log-likelihood $\log p_{\theta}(x)$ can be obtained according to the change of variables formula:
$\begin{split}\log p_{\theta}(x)&=\log p_{Z}(z)+\log|\det(\frac{\partial f_{\theta}}{\partial x})|\\ &=\log p_{Z}(z)+\log|\det\left(J_{f_{\theta}}(f_{\theta}^{-1}(z))\right)|,\end{split}$ (1)
where $J_{f_{\theta}}(\cdot)$ calculates the Jacobian matrix of the transformation $f_{\theta}(x)$, and $\det(\cdot)$ computes the matrix determinant.
It should be noted that $f_{\theta}$ can be composed of multiple layers in the case of using deep learning models, and each layer needs to be an invertible function to guarantee the invertibility of $f_{\theta}$:
$\begin{split}f_{\theta}(x)&=f_{L}\circ f_{L-1}\circ\dots\circ f_{1}(x),\\ f_{\theta}^{-1}(z)&=f_{1}^{-1}\circ f_{2}^{-1}\circ\dots\circ f_{L}^{-1}(z),\end{split}$ (2)
where $\circ$ is for function composition and there are $L$ layers in the model. According to the chain rule in differentiability, the determinant of $f_{\theta}$ is now transformed into layer-wise computation:
$\log\left|\frac{\partial f_{\theta}}{\partial x}\right|=\log\left|\frac{\partial f_{1}}{\partial x}\right|+\sum_{i=2}^{L}\log\left|\frac{\partial f_{i}}{\partial f_{i-1}}\right|.$ (3)
Therefore, it is important to have an efficient log-determinant computation when training normalizing flow models. When $f_{i}$ is a linear layer with weight matrix $\mathbf{W}$, its Jabobian matrix is equivalent to the weight matrix:
$\log|\det\frac{\partial f_{i}}{\partial f_{i-1}}|=\log|\det\mathbf{W}|,$ (4)
where the determinant computation complexity is $\mathcal{O}(n^{3})$.
Normalizing flow model can perform data generation using $f^{-1}_{\theta}$ by feeding with samples $z$ from $p_{Z}(z)$. Therefore, it is important to have an efficient $f^{-1}_{\theta}$ to achieve fast data generation. In particular, when $f_{i}$ is a linear layer, its inverse is equivalent to the linear transformation built on top of the inverse of weight matrix $\mathbf{W}$:
$f_{i}^{-1}(x)=\mathbf{W}^{-1}\times x.$ (5)
In general, the matrix inverse can take $\mathcal{O}(n^{3})$ computation complexity.
Overall, log-determinant computation needs to be fast for training normalizing flow models, while data generation requires fast inverse computation.
Nonlinear Coupling Layer.
The coupling layer (Dinh et al., 2014, 2016) is a simple and effective invertible transformation, where the Jacobi matrix is a triangular matrix. It divides the input into two parts $x_{a}$ and $x_{b}$, then transforms $x_{b}$ using a scaling parameter $s_{\theta}$ and a translation parameter $b_{\theta}$ that can be a nonlinear function of $x_{a}$, while $z_{a}$ is directly equal to $x_{a}$, and the final output concatenates the two parts together:
$z_{a}=x_{a},$ (6)
$z_{b}=s_{\theta}\odot x_{b}+b_{\theta},$ (7)
$z=\text{concat}(z_{a},z_{b}).$ (8)
Coupling layers remain a crucial component in most flow-based models (Kingma and Dhariwal, 2018; Hoogeboom et al., 2019; Meng et al., 2022). However, since coupling layers only transform a subset of the variables, they must be combined with linear layers to facilitate information fusion across channels. As a result, recent research has increasingly focused on the design of linear layers to enhance model performance.
Invertible Linear Layer.
A well-designed linear layer can significantly enhance both the coupling layer’s ability to learn the target distribution and the overall expressive power of the model. Furthermore, the computational cost associated with matrix inversion and determinant computation serves as an important indicator of the efficiency of linear layer design.
Glow (Kingma and Dhariwal, 2018) pioneered the use of $1\times 1$ convolution with learnable parameters as a linear layer in flow-based models. This method is more effective at capturing correlations between variables compared to simpler channel shuffling operations (Dinh et al., 2014, 2016). Additionally, Glow introduces a weight design based on LU decomposition, reducing the Jacobian determinant computation from $\mathcal{O}(n^{3})$ to $\mathcal{O}(n)$.
Similarly, (Hoogeboom et al., 2019) extended convolution to $d\times d$ and proposed periodic convolution, leveraging the Fast Fourier Transform (FFT) to accelerate the convolution process. However, the periodic convolution method involves generalized determinant computation, leading to a complexity of $\mathcal{O}(n^{3})$. In contrast, our proposed method not only accelerates the convolution process using FFT, but also reduces the determinant computation complexity to $\mathcal{O}(n)$.
Despite numerous works exploiting special matrix properties to enhance model expressiveness, most focus on improving the efficiency of the Jacobian determinant computation, with limited optimization of the matrix inversion process during the sampling phase of the flow model. In these methods, including the aforementioned works (Kingma and Dhariwal, 2018; Hoogeboom et al., 2019; Meng et al., 2022), the computational complexity for the inverse sampling operation is typically $\mathcal{O}(n^{2})$. The Woodbury Transformation (Lu and Huang, 2020) achieves a theoretical complexity of $\mathcal{O}(dn)$, where $d$ is the dimensionality of smaller matrices (e.g., 8 or 16). Although this is better in theory than the $\mathcal{O}(n\log n)$ complexity of our method, in practice Woodbury requires multiple input transformations, which makes its actual efficiency inferior to ours.
2.2 Weight Matrix Construction and Fast Multiplication
A general matrix $\mathbf{M}\in\mathbb{R}^{n\times n}$ can be factorized into the alternating product of circulant and diagonal matrices, with the total number of matrices not exceeding 2n12n-1 (Huhtanen and Perämäki, 2015). Specifically, it can be written as:
$\mathbf{M}=\mathbf{D}_{1}\mathbf{C}_{2}\mathbf{D}_{3}\dots\mathbf{D}_{2n-3}\mathbf{C}_{2n-2}\mathbf{D}_{2n-1},$ (9)
where $\mathbf{D}_{2j-1}$ and $\mathbf{C}_{2j}$ are diagonal and circulant matrix for $j=1,2,\dots,n$. This decomposition can theoretically approximate any dense matrix, and the approximation error can be controlled by adjusting the number of factors. (Ding et al., 2025) have successfully applied such representation to fine-tuning large language models.
Denote the input vector as $\mathbf{x}\in\mathbb{R}^{n}$. The weight matrix $\mathbf{W}\in\mathbb{R}^{n\times n}$ can be represented by $m$ diagonal matrices and m1m-1 circulant matrices, thereby $m\leq n$. For each $j\in\{1,2,\dots,m\}$, the $j$-th diagonal matrix is defined by the vector $\mathbf{d}_{2j-1}\in\mathbb{R}^{n\times 1}$, and the $j$-th circulant matrix is defined by the vector $\mathbf{c}_{2j}\in\mathbb{R}^{n\times 1}$. Thus, the model weight matrix can be represented as:
$\mathbf{W}$ $=\mathbf{D}_{1}\mathbf{C}_{2}\mathbf{D}_{3}\dots\mathbf{D}_{2m-3}\mathbf{C}_{2m-2}\mathbf{D}_{2m-1}$
$=\text{diag}(\mathbf{d}_{1})\times\text{circ}(\mathbf{c}_{2})\times\dots\times\text{diag}(\mathbf{d}_{2m-1}),$ (10)
where $\text{diag}(\cdot)$ and $\text{circ}(\cdot)$ represent the construction of diagonal and circulant matrices, respectively. It should also be noted that circulant matrix can be factorized into:
$\text{circ}(\mathbf{c}_{2j})=\mathbf{F}^{-1}\times\text{diag}(\hat{\mathbf{c}}_{2j})\times\mathbf{F},\quad\hat{\mathbf{c}}_{2j}=\mathbf{F}\times\mathbf{c}_{2j},$ (11)
where $\mathbf{F}$ is the order $n$ discrete Fourier transform matrix. In practice, it is better to implement with the frequency domain parameter $\hat{\mathbf{c}}_{2j}$ as weight parameter rather than real valued $\mathbf{c}_{2j}$ to save computation costs in matrix vector product, logarithm determinant, and matrix inversion.
3 Invertible Linear Layer Design
In this section, we introduce our invertible linear layer that consist of circulant and diagonal matrices. We discuss the fast algorithms for the matrix vector product, logarithm of matrix determinant and matrix inversion. Additionally, we also analyze the theoretical computation cost associated with these operations.
3.1 Matrix Log-determinant
According to Eq. 10, The weight matrix is constructed as an alternating product of diagonal and circulant matrices. The determinant of this matrix can be efficiently computed by leveraging the special properties of these matrices. Recall that determinant of matrix product is equal to product of matrix determinant:
$\det(\mathbf{W})$ $=\prod_{j=1}^{m}\det(\mathbf{D}_{2j-1})\times\prod_{j=1}^{m-1}\det(\mathbf{C}_{2j}),$ (12)
$\det(\mathbf{D}_{2j-1})$ $=\det(\text{diag}(\mathbf{d}_{2j-1}))=\prod_{i=1}^{n}d_{2j-1}^{i},$ (13)
$\det(\mathbf{C}_{2j})$ $=\det(\mathbf{F}^{-1})\det(\text{diag}(\hat{\mathbf{c}}_{2j}))\det(\mathbf{F})=\prod_{i=1}^{n}\hat{c}_{2j}^{i}.$ (14)
To compute the logarithm of the matrix determinant of $\mathbf{W}$, we can express into the sum of each matrix determinant:
$\log|\det(\mathbf{W})|=\sum_{j=1}^{m}\sum_{i=1}^{n}\log|\hat{c}_{2j}^{i}|+\sum_{j=1}^{m-1}\sum_{i=1}^{n}\log|d_{2j-1}^{i}|.$ (15)
Therefore, the overall log-determinant computation can be finished by simple summation with a complexity of $\mathcal{O}(mn)$.
3.2 Matrix Inverse
The inversion of the weight matrix for the data sampling process can be efficiently implemented by leveraging the decomposition of the weight matrix into alternating diagonal and circulant matrices. This section introduces the inverse calculation process, including its objective, step-by-step procedure, and computational complexity. The objective of the matrix inversion is to recover the input vector $\mathbf{x}$ from the output $\mathbf{y}$, given the weight matrix $\mathbf{W}$. This is achieved by solving the equation:
$\mathbf{x}=\mathbf{W}^{-1}\times\mathbf{y},$ (16)
where the inverse of $\mathbf{W}$ is represented as:
$\mathbf{W}^{-1}=\mathbf{D}_{2m-1}^{-1}\times\dots\times\mathbf{C}_{2}^{-1}\mathbf{D}_{1}^{-1}.$ (17)
Diagonal Matrices.
For each diagonal matrix $\mathbf{D}_{2j-1}$, its inverse is obtained by taking the reciprocal of its diagonal elements.
$\mathbf{D}_{2j-1}^{-1}=\text{diag}(1/d_{2j-1}^{1},1/d_{2j-1}^{2},\dots,1/d_{2j-1}^{n}),$ (18)
which has a computation complexity of $\mathcal{O}(n)$.
Circulant Matrices.
For each circulant matrix $\mathbf{C}_{j}$, its inverse is similar to a diagonal matrix according to Eq. (11):
$\begin{split}\mathbf{C}_{2j}^{-1}&=\mathbf{F}^{-1}\times\text{diag}^{-1}(\hat{\mathbf{c}}_{2j})\times\mathbf{F}\\ &=\mathbf{F}^{-1}\times\text{diag}(1/\hat{c}^{1}_{2j},\dots,1/\hat{c}^{n}_{2j})\times\mathbf{F}.\end{split}$ (19)
It can be seen that matrix inversion $\mathbf{W}^{-1}$ can be immediately determined given $\mathbf{d}_{2j-1}$ and $\mathbf{c}_{2j}$. Therefore, the complexity of Eq. (16) is equivalent to matrix vector product complexity, i.e., $\mathcal{O}(mn\log n)$.
Table 1: Comparisons of computational complexities and parameter counts, where $n$ is the number of input channels. In practice, $d$ is the dimensionality of the smaller matrices used in the Woodbury transformation (typically $d=8$ or 1616), $L$ is the number of butterfly layers, and $m$ is generally set to 2 for our method. Here $h$ and $w$ denote the spatial dimensions of the feature maps.
Method Logdet Inverse Params
$1\times 1$ Conv (Kingma and Dhariwal, 2018) $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$
$1\times 1$ Conv (LU) (Kingma and Dhariwal, 2018) $\mathcal{O}(n)$ $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$
Emerging (Hoogeboom et al., 2019) $\mathcal{O}(n)$ $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$
Periodic (Hoogeboom et al., 2019) $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$
Woodbury (Lu and Huang, 2020) $\mathcal{O}(nd)$ $\mathcal{O}(nd)$ $\mathcal{O}(d(n+h\cdot w))$
ME-Woodbury (Lu and Huang, 2020) $\mathcal{O}(nd)$ $\mathcal{O}(nd)$ $\mathcal{O}(d(n+h+w))$
ButterflyFlow (Meng et al., 2022) $\mathcal{O}(n)$ $\mathcal{O}(n^{2})$ $\mathcal{O}(nL)$
CDFlow (Ours) $\mathcal{O}(mn)$ $\mathcal{O}(mn\log n)$ $\mathcal{O}(mn)$
3.3 Generative Modeling with CDFlow
Figure 1: Architecture of the CDFlow Model. The left portion shows a detailed schematic of the proposed CD-Convolution layer, which is constructed using circulant and diagonal matrices. The middle portion illustrates the internal structure of a flow module, where the input sequentially passes through an Actnorm layer(Kingma and Dhariwal, 2018), the CD Convolution layer, and a coupling layer(Dinh et al., 2016). The right portion presents the overall CDFlow model architecture, in which the flow module is repeated K times and integrated into a multi-scale framework(Dinh et al., 2016) to effectively fuse channel-wise information.
We present our CDFlow model that consists of $L$ blocks, where each block is formed by stacking $K$ layers of flow modules. Each flow module consists of the ActNorm layer, the invertible convolutional layer, and the coupling layer, working together to transform the input data. The multi-scale structure helps the model learn complex relationships within the data by processing it at different levels of abstraction. Figure 1 illustrates how the flow module is constructed for each layer and how the entire model is repeated and integrated to learn and generate the target data.
Actnorm Layers.
The ActNorm layer (Kingma and Dhariwal, 2018) is a type of normalization technique that uses affine transformations to perform class normalization. Unlike traditional methods like BatchNorm (Ioffe and Szegedy, 2015), which rely on the statistics of the entire batch, the ActNorm layer initializes its parameters based on the statistics of the first batch. Once initialized, these parameters are treated as learnable variables. This initialization method is advantageous for small batch sizes, as it avoids the performance degradation that BatchNorm often encounters when dealing with small batches. ActNorm helps in stabilizing the training of normalizing flows and is especially useful for flow-based models that require flexible scaling and translation invariance.
Invertible Convolutional Layer.
In CDFlow, the CD-Convolution layer leverages special weight matrices, specifically circulant and diagonal matrices, to enable efficient computation. This design significantly reduces the computational cost of matrix inversion and the computation of the Jacobi determinant, which are essential operations for flow-based models. A current limitation of our approach is its reliance on $1\times 1$ convolution, and our future work will extend to $d\times d$ convolutions. A detailed explanation of this design can be found in Section 3.
Coupling Layers.
The coupling layer (Dinh et al., 2014, 2016) is a core component of normalizing flow models. The coupling layer splits the input into two parts: one part is passed through a neural network to compute a transformation, while the other part is left unchanged. This operation enables the model to learn complex transformations while maintaining a simple and computationally efficient structure. The coupling layer is combined with the reversible convolutional layer to form the flow module, allowing the CDFlow model to capture complex distributions and generate high-quality samples.
Multi-scale architecture.
CDFlow incorporates a multi-scale architecture (Dinh et al., 2016), which improves the model’s ability to capture complex and high-dimensional data distributions. Within each block, the split and squeeze operations play an important role:
  • The split operation divides the input tensor along the channel dimension, which allows the model to independently process different features. This enhances the model’s expressive power by enabling it to focus on diverse aspects of the data.
  • The squeeze operation reduces the spatial resolution of the data by downsampling the spatial dimensions. This operation allows the model to focus on higher-level features while also reducing computational cost by focusing on the most important data representations.
4 Experiments
In this section, we present the experimental results to evaluate the performance of the proposed method. The experiments are conducted in three main parts:
  1. 1.
    We first evaluate the method on commonly used natural image datasets to demonstrate its general effectiveness under standard conditions.
  2. 2.
    Next, we apply the method to datasets with specialized structures to investigate its performance on data exhibiting periodic patterns.
  3. 3.
    Finally, we conduct a runtime analysis to assess the computational efficiency of the method.
The following subsections present the detailed results and discussions for each of these experiments.
We further investigated the effectiveness of the proposed linear layer within the Flow Matching model, and the results can be found in Appendix  C.
4.1 Experimental Settings and Model Parameters
Datasets and Baselines.
To compare our method with previous approaches, we evaluate its performance on the CIFAR-10 (Krizhevsky, 2009) and ImageNet  (Deng et al., 2009) datasets, using bits per dimension (BPD) as the evaluation metric. BPD is a commonly used measure for assessing the quality of generative models, indicating the number of bits required to encode each dimension of the data. Metrics such as IS and FID are not used in this work, as they do not generalize well across different datasets and fail to account for overfitting. For the baseline, we select several models related to CDFlow, including MAF (Papamakarios et al., 2017), Real NVP (Dinh et al., 2016), Glow (Kingma and Dhariwal, 2018), Emerging (Hoogeboom et al., 2019), i-ResNet (Behrmann et al., 2019), Woodbury (Lu and Huang, 2020) and ButterflyFlow (Meng et al., 2022).
Implementation Details.
To facilitate a fair comparison with prior architectures, we adopt consistent backbone settings for flow-based baselines that share a similar multi-scale design, including Glow, Woodbury, Emerging, and ButterflyFlow. On standard images datasets, all of these models are configured with three blocks and 32 flow steps. On the structured dataset, we instead adopt a lighter configuration with two blocks and eight flow steps. For Woodbury, we adhere to the original configuration with $d_{s}=d_{c}=16$, while for ButterflyFlow we use the default setting with butterfly level =1=1. For our proposed CDFlow, we set $m=2$, meaning we use two diagonal vectors and a circulant vector. The experiments were executed on NVIDIA A100 and RTX 4090 GPUs.
During training, we apply spectral normalization (Miyato et al., 2018) to the convolutional layers to ensure stability. Thanks to the structure of circulant and diagonal matrices, the maximum singular value of each weight matrix can be efficiently computed and used to rescale the weights. In addition, we employ a channel-aware learning rate scaling for the CD-Convolution parameters, which prevents abrupt updates to the structured matrices. Empirically, this combination further stabilizes the training of CDFlow.
Figure 2: Ablation study on the hyperparameter $m$. Left: parameter counts (K) of the CD-Convolution layers and BPD on CIFAR-10. Middle: forward computation time (ms) across different channel sizes. Right: inverse computation time (ms) across different channel sizes.
Hyperparameter Selection
Each linear transform in our method is decomposed into 2m12m-1 matrices (i.e., $m$ diagonal and m1m-1 circulant). While larger $m$ values increase the approximation power, they also introduce additional parameter and computational costs. The key objective is to balance expressiveness, efficiency, and runtime performance. To this end, we systematically evaluate the effect of varying $m$ on model behavior.
As shown in Figure 2(a), increasing $m$ on CIFAR-10 leads to only marginal improvements in expressiveness, as reflected by BPD values, while significantly increasing the parameter count of the CD-Convolution layers. Figures 2(b) and (c) further report the runtime measurements for forward and inverse operations. Although larger $m$ values introduce additional latency, our method remains more efficient than the other methods reported in Section 4.4. In addition, we separately evaluated the log-determinant computation time across channel sizes ranging from 16 to 1024, where the runtime remains consistent across different $m$, with values between 0.095 ms and 0.097 ms.
Based on these results, we adopt $m=2$ as the default setting for all experiments, including both the main and supplementary studies, unless otherwise specified.
4.2 Density Estimation on Standard Images
Standard Image Datasets.
To enable direct comparison with prior work (Meng et al., 2022), we evaluate our method on the CIFAR-10 (Krizhevsky, 2009) and ImageNet-32×32 (Deng et al., 2009) datasets. For both datasets, training is conducted with unified dequantization and standard data augmentation techniques.
Results.
Table 2 presents the quantization results of the model on the CIFAR-10 and ImageNet $32\times 32$ datasets. CDFlow outperforms Glow, Emerging, Woodbury, and ButterflyFlow (all using the same modeling framework) and performs comparably to models like Residual Flows and i-DenseNet. The generated sample results are visualized in Figure 3(a) and Figure 3(b).
Table 2: Density estimation results on CIFAR-10, ImageNet $32\times 32$, and Galaxy, including the corresponding model sizes (Params) and bits per dimension (BPD). Lower BPD indicates better performance.
Model CIFAR-10 ImageNet $32\times 32$ CIFAR-100
Params BPD ($\downarrow$) Params BPD ($\downarrow$) Params BPD ($\downarrow$)
MAF (Papamakarios et al., 2017) 105.0M 4.31
Real NVP (Dinh et al., 2016) 6.4M 3.49 46.2M 4.28 6.38M 2.11
Glow (Kingma and Dhariwal, 2018) 44.2M $3.36\pm 0.002$ 66.2M 4.09 6.25M 2.02
Emerging (Hoogeboom et al., 2019) 46.6M $3.34\pm 0.002$ 44.0M 4.09 6.68M 1.98
Residual Flows (Chen et al., 2019) 25.2M 3.28 47.1M 4.01 5.58M 3.60
i-ResNet (Behrmann et al., 2019) 25.0M $3.32\pm 0.006$ 29.7M 4.05 6.88M 2.04
Woodbury (Lu and Huang, 2020) 45.3M $3.42\pm 0.002$ 45.3M 4.09 6.68M 2.01
i-DenseNet (Perugachi-Diaz et al., 2021) 24.9M 3.25 47.0M 3.98 6.07M 4.06
ButterflyFlow (Meng et al., 2022) 44.4M $3.33\pm 0.003$ 44.4M 4.09 6.39M 1.95
CDFlow (Ours) 44.2M $3.31\pm 0.004$ 44.2M 4.04 6.25M 1.92
(a) CIFAR-10
(b) ImageNet $32\times 32$
(c) Galaxy
Figure 3: Comparison between the original image and outputs from different methods. For clearer and larger visualizations, please refer to Appendix D.
4.3 Density Estimation on Structured Images
To demonstrate the capability of CDFlow in modeling periodic structures, we apply it to galaxy images (Ackermann et al., 2018), conducting an experiment on this real-world dataset.
Galaxy Datasets.
When the boundary pixels of an image are relatively uniform or connected, the data can be considered periodic. A typical example of periodic data is galaxy images, which often feature scattered light sources with predominantly dark boundaries. The Galaxy datasets (Ackermann et al., 2018) is a small classification dataset that includes 3,000 merged galaxies and 5,000 non-merged galaxies in the training and test sets. We train CDFlow on non-merged galaxy images and compare its performance to other models under the same experimental setup.
Results.
Table 2 also presents the BPD results on the galaxy dataset. It is clear that CDFlow outperforms all other models, achieving a 4.95% improvement in BPD compared to the Glow model. In addition, Figure 3(c) presents a visual comparison between our method and two strong baselines: Emerging and Butterfly. Experimental results demonstrate that our proposed CDFlow model can well model data with periodic qualities.
4.4 Runtime
To demonstrate the effectiveness of our method in reducing computational complexity, we separately evaluated the runtime of the forward operation, log-determinant (logdet) computation, and inverse operation. We compared our method with standard $1\times 1$ convolution, emerging and periodic convolutions (Hoogeboom et al., 2019), the Woodbury transformation and ME-Woodbury with a fixed latent dimension of $d=16$ (Lu and Huang, 2020), and the butterfly layer with level 10 (Meng et al., 2022). For a fair comparison, all methods were implemented in PyTorch, and all experiments were conducted on an NVIDIA A800 GPU.
Figure 4: Comparison of runtime (ms) for forward computation, inverse computation, and log-determinant calculation under varying numbers of channels, as well as inverse computation under different spatial sizes. The results represent the average of 100 runs. Our method consistently outperforms baselines, especially in inverse and log-determinant computations. The runtime of the Emerging method becomes excessively long at large spatial sizes and is therefore omitted from the corresponding figure.
Forward Time.
The forward operation refers to the runtime required to apply the tested layer function and compute the log-determinant, i.e., $y=f(x)$ and $\log|\det(\mathbf{W})|$, under gradient computation mode. As shown in Figure 4, although the theoretical complexity of the $1\times 1$ convolution is $\mathcal{O}(n^{3})$, its actual runtime remains relatively stable across varying channel sizes, likely due to GPU parallelism. All baseline methods are slower than the $1\times 1$ convolution, while our method achieves even faster performance. Notably, although the periodic convolution leverages Fourier transforms, its reliance on two-dimensional FFT introduces substantial memory overhead, leading to sharp runtime increases with more channels. In contrast, our method combines circulant and diagonal matrix constructions to reduce log-determinant complexity, and employs an optimized convolutional implementation to accelerate forward computation, as detailed in Sections 2.2 and 3.1.
In addition, we specifically measured the runtime of the log-determinant computation, which depends solely on the size of the weight matrix, i.e., the number of input channels. Figure 4 illustrates how the runtime of different methods varies with increasing channel dimensions. Our method not only outperforms all baselines but also maintains stable performance without significant growth as the number of channels increases.
Inverse Time.
The inverse operation refers to the time required to apply the inverse function of the tested layer, i.e., $x=f^{-1}(y)$, under no-gradient mode. We first analyze the impact of channel number on runtime, as shown in Figure 4. The Emerging convolution suffers from poor parallelism, while the periodic convolution relies on two-dimensional Fourier transforms, leading to a significant increase in runtime as the number of channels grows. Other baseline methods perform slightly worse than the $1\times 1$ convolution, possibly due to the lack of optimized inverse implementations and additional transformation overhead. In contrast, our method reduces the inversion complexity to $\mathcal{O}(n\log n)$, and despite requiring data reshaping, consistently outperforms other baselines, particularly at large channel sizes.
We further evaluate how spatial resolution affects runtime, with the number of channels and batch size fixed. The results show that all methods experience increased runtime with larger spatial sizes. Our method outperforms the $1\times 1$ convolution at small spatial sizes, and becomes slightly less efficient at larger sizes, since it mainly reduces channel-wise rather than spatial complexity. Nevertheless, our method still clearly outperforms all other baselines except the $1\times 1$ convolution.
5 Conclusion
In this paper, we introduce the CDFlow model, a novel flow-based generative model. Instead of relying on traditional weight matrices, we employ the interleaved multiplication of circulant and diagonal matrices, leveraging the properties of these matrices along with the Fast Fourier Transform (FFT) to efficiently compute the Jacobian determinant and perform the inverse transformation. Our experiments demonstrate that CDFlow excels in density estimation for standard image datasets and performs well on real-world data with periodic properties, outperforming existing baselines.
Acknowledgments and Disclosure of Funding
This work was financially supported by the National Key R&D Program of China (Grant No. 2024YFA1211400).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: As described in the abstract and introduction, our paper proposes a new normalization flow model that accelerates matrix inversion and determinant computation through weight decomposition and Fourier transform.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Sections 3.3 and 4.4 discuss limitations, including the current restriction to $1\times 1$ convolutions and reduced efficiency on large spatial inputs.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Section 3 presents the mathematical derivation of our approach, including the assumptions and computation process for the transformation and determinant.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Section 4 and the appendix provides sufficient experimental details to allow reproduction, including dataset descriptions, training procedures, and model configurations.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We plan to release the anonymized version of the code and data upon publication.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Section 4.1 and the appendix provides describes training and testing details.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We do not report error bars, but we provide multiple comparative experiments to support the main claims.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Section 4.1 and Section 4.4 include information on the GPU model used and the runtime, offering guidance for reproduction in similar environments.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: This work does not involve any human subjects, sensitive data, or potential risks; it fully conforms with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [No]
  49. Justification: We did not explicitly discuss broader impacts, as this work focuses on a theoretical contribution to machine learning model efficiency, without direct deployment or user interaction components.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our work does not involve the release of models or datasets with high risk of misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All external datasets and baseline models used are cited properly in the main text.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should citet the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We will provide documentation and usage instructions for our code and model assets in the supplemental material upon release.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing or human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This work does not involve human subjects, hence IRB approval is not required.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLMs were only used for language editing and polishing purposes.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Impact of Linear Transformation Constraints on Flow-based Model Performance
Figure 5: NLL comparison among flow-based models constructed with different linear layer types (F, L, U, LU, and DCD). Our DCD-type layer achieves performance close to the fully learnable F-type and surpasses the LU-type structure.
To further validate the effectiveness of our proposed method, we reproduced the experiment described in Appendix A.5 of Chao et al. [2023] and included an additional variant based on our Diagonal-Circulant Decomposition (DCD) linear layers. The compared linear layer types include F-type (full matrix), L-type (lower-triangular), U-type (upper-triangular), LU-type (the product of lower and upper triangular matrices), and our proposed DCD-type.
As shown in Fig. 5, the model employing the F-type layers achieved the lowest negative log-likelihood (NLL), consistent with Chao et al. [2023]. Although our DCD-type layer does not surpass the fully learnable F-type, it achieves the second-best performance, outperforming the LU-type design. This result indicates that DCD retains much of the expressiveness of the unconstrained full matrix while maintaining a more efficient and structured parameterization.
Appendix B Runtime
Specifically, our approach requires transforming the input tensor from a four-dimensional structure $[B,C,H,W]$ to a two-dimensional matrix $B\times H\times W,C]$. While the runtime reported in Section 4.4 omits this transformation, we further measured the total runtime including it to ensure a fair comparison. As illustrated in Figure 6, our method continues to demonstrate strong computational efficiency even after accounting for the transformation overhead. It is worth noting that the computation of the log-determinant is independent of the input tensor, and therefore the reshaping operation has no impact on its runtime.
Figure 6: Comparison of runtime (ms) .
Appendix C Additional Experiments on Flow Matching Models
To further validate the effectiveness of the proposed linear layer, we conducted experiments within the Flow Matching framework Lipman et al. [2022], which employs a 5-layer MLP on 2D toy datasets. To systematically evaluate the method, we performed experiments on three datasets: Chessboard, Moons, and Circles. Within the original framework, we replaced the linear layers of the standard MLP with four structured variants for comparison: EMLP (Emerging Layer) Hoogeboom et al. [2019], WMLP (Woodbury Layer) Lu and Huang [2020], BFMLP (Butterfly Layer) Meng et al. [2022], and CDMLP (Ours).
The experimental results are summarized in Table 3. CDMLP achieves the best or near-best performance across most cases while requiring significantly fewer parameters. For instance, it attains the lowest NLL on the Chessboard and Circles datasets and the best MSE and FID on Moons, outperforming or matching larger models such as MLP and EMLP. In addition, Figures 78 further visualize the learned distributions, showing that the model successfully evolves from Gaussian initialization toward the target distributions. These findings not only confirm the modeling capability of the proposed method within the Flow Matching framework but also highlight its applicability and generality across different data structures.
Table 3: Flow Matching Model Results on 2D-toy Datasets. The best results are highlighted in bold.
[…]
Chessboard MLP Lipman et al. [2022] 791.0K 3.796 0.0062 10.6853
EMLP Hoogeboom et al. [2019] 1577.5K 3.797 0.0165 9.357
WMLP Lu and Huang [2020] 40.5K 3.761 0.0042 9.0551
BFMLP Meng et al. [2022] 26.1K 3.789 0.0125 8.2088
CDMLP (Ours) 8.45K 3.810 0.0115 7.0425
Moons MLP Lipman et al. [2022] 791.0K 1.854 0.0022 24.1885
EMLP Hoogeboom et al. [2019] 1577.5K 1.855 0.0023 36.9462
WMLP Lu and Huang [2020] 40.5K 1.905 0.0016 47.6328
BFMLP Meng et al. [2022] 26.1K 1.886 0.0031 17.9168
CDMLP (Ours) 8.45K 1.842 0.0016 31.0329
Circles MLP Lipman et al. [2022] 791.0K 2.433 0.0019 42.4687
EMLP Hoogeboom et al. [2019] 1577.5K 2.517 0.0032 20.5252
WMLP Lu and Huang [2020] 40.5K 2.547 0.0021 26.6795
BFMLP Meng et al. [2022] 26.1K 2.470 0.0032 21.6716
CDMLP (Ours) 8.45K 2.482 0.0069 15.4922
Figure 7: Time evolution of the learned distribution on 2D toy datasets (Chessboard, Moons, Circles).
Figure 8: Visualization of model likelihood (left: Hutchinson estimator, right: exact) on 2D toy datasets (Chessboard, Moons, Circles).
Appendix D Visualization of Generation Results on Image Datasets
Figure 9: Generation results on the CIFAR-10 dataset.
Figure 10: Generation results on the ImageNet dataset.
Figure 11: Generation results on the Galaxy dataset.